# sections/MRM-supplementary

### v1

@@ 4 Discussion
In addition to faster processing times, WALINET showed superior performance with more lipid removal and preserving more metabolite signal compared to state-of-the-art conventional lipid removal methods. WALINET provided significantly higher data quality, effectively doubling the SNR and lowering by half quantification errors of metabolites, compared to conventional methods.
Convolutional neural networks act as nonlinear functions that may model better lipid contamination in MRSI, while conventional methods 29, 23 that assume a linear orthogonal relationship between lipids and metabolites may inadvertently remove metabolite signal when this assumption is not met. The effects of improved metabolite quantification translate into metabolite images that have better structural details. We believe that a robust self-contained efficient nuisance signal removal method that is implemented as an independent processing step is very useful and can be combined with any MRSI processing pipeline. This may offer greater flexibility compared to methods that are fully embedded with reconstruction of k-space MRSI data.
@@
At the moment, the demonstration of WALINET performance was limited to 7T ultra-high-field MRSI, which is the highest field approved for clinical use. Methods such as WALINET are highly relevant at 7T because the high SAR and non-uniform $B_{0}$ and $B_{1}+$ fields make pulse sequence-based suppression of water and lipids highly impractical for whole-brain MRSI. However, we expect that the approach and same network can be employed with additional training at lower (3T) or higher fields. We also expect that the processing time of WALINET will be similar for higher spatial resolution of MRSI, while the time for conventional methods will linearly scale with the data size. Furthermore, removal of lipid signal is extremely important for accelerated undersampled MRSI acquisitions 48, 49 where aliased lipid signal can overwhelm parallel imaging and compressed sense reconstructions. Based on its performance, WALINET has great potential for being used in combination with undersampled MRSI, and we will explore this in future work. This is suggested by the generalization of WALINET from 3D Non-Cartesian sampling to 2D Cartesian sampling that we demonstrated in this paper. In addition, we showed that the same model can be used for lipid-only removal (LIPNET), which can be further combined with other processing methods. We also noticed that LIPNET provides some suppression of the water tail in the metabolite spectral range. This happens because there are lipid signals (at 5.1ppm and 4.4 ppm) close to the water peak which the network learns to remove. However, in the presence of a large residual water peak LIPNET is not sufficient and a water trained network is necessary such as WALINET.
@@ Conclusion
Efficient removal of water and lipid signals is key for proton MRSI-based quantitative metabolic imaging. Convolutional neural networks such as WALINET provide an effective approach towards this goal with superior performance compared to conventional methods. We provide WALINET, including the computation of the L2 lipid operator, as a self-contained package that can be used as a plug-in with any MRSI processing pipeline. WALINET can be used to automate the MRSI processing pipeline for high-throughput workflow in clinical applications. We anticipate that these aspects will lead to larger adoption and impact of MRSI for clinical applications and research.
@@ 5 Supplementary Methods
2D 1H-FID Cartesian MRSI data were acquired with unaccelerated elliptical phase-encoded. 3D 1H-FID-ECCENTRIC 46 was acquired with randomly positioned circular trajectories with the radius set to kmax/8 without temporal interleaving and full sampling (AF=1) of a spherical 3D k-space.
For both sequences the excitation was performed with a Shinnar-LeRoux optimized pulse 45, 46 having 6.5kHz bandwidth, 1ms duration and 27° excitation flip-angle. No lipid suppression was employed in the sequence, while water suppression was achieved by a four-pulses WET method 45, 46.
@@
For 3D ECCENTRIC the k-space sampling density was compensated based on Voronoi diagrams 50 where each k-space point is normalized by the area of its assigned Voronoi vertex. Upon the weighting of the k-space data, an inverse non-uniform discrete Fourier transform (iNUFT) was applied for each stack (kx-ky plane) within the k-space domain. Subsequently, an additional inverse fast Fourier transform (iFFT) was performed along the kz dimension to finalize the reconstruction of the MRSI data. In contrast, for 2D ${}^{1}H$-FID-MRSI data a iFFT is applied on the 2D k-space datasets. For both 2D and 3D MRSI a Hamming filter was applied in k-space prior to the Fourier transform. After the transformation from k space to image space, the spectra are obtained by Fourier transform of the time dimension. Coil combination was performed with ESPIRIT 51 using sensitivity profiles computed from water un-suppressed MRSI. A correction for $B_{0}$ field inhomogeneity (Supp. Fig. 1) is computed from water un-suppressed data and applied to the coil-combined image space MRSI data.
@@
LCModel used a basis set simulated by NMR quantum mechanics in GAMMA 52 for twenty-two metabolites: phosphorylcholine (PCh), glycerophosphorylcholine (GPC), creatine (Cr), phosphocreatine (PCr), gamma-aminobutyric acid (GABA), glutamate (Glu), glutamine (Gln), glycine (Gly), glutathione (GSH), myo-inositol (Ins), N-acetylaspartate (NAA), N-acetyl aspartylglutamate (NAAG), scylloinositol (Sci), lactate (Lac), threonine (Thr), beta-glucose (bGlu), alanine (Ala), aspartate (Asp), ascorbate (Asc), serine (Ser), taurine (Tau), and 2-hydroxyglutarate (2HG) and a measured macromolecular background 53. Note that during training WALINET and LIPNET learn to remove macromolecule signal, since this is present together with lipid signal in the scalp spectra used to generate training data. This was verified experimentally, as macromolecular fitting by LCModel was very close to 0 throughout the brain. The spectral fitting was done for the 1ppm-4.2ppm spectral range and the results for each voxel were used to generate metabolic images. The unsupressed water reference signal was used as quantification reference for metabolites concentrations (institutional units, I.U.) to compare metabolite levels across subjects and scanners. To assess the quality of the MRSI data and fit, linewidth (FWHM), signal-to-noise ratio (SNR), and Cramer-Rao lower bounds (CRLB) goodness of fit maps were generated.
@@
Supplementary Figure 1: Examples of head/brain/scalp binary masks and the corresponding B0 field maps (specified in ppm).
@@
Supplementary Figure 2: Comparison of lipid removal on in-vivo 2D MRSI by LIPNET and L2 for three values of the regularization parameter $\beta$, the optimal value ($3.69*10^{6}$), double the optimal value, and zero for no lipid suppression. Maps are shown for two metabolites Cr+PCr, Glu and their corresponding CLRB. Spectra from several brain voxels are shown for each method, the white trace shows the measured spectrum, the red trace shows LCModel fit.

### v2

@@ 4 Discussion
In addition to faster processing times, WALINET showed superior performance with more lipid removal and preserving more metabolite signal compared to state-of-the-art conventional lipid removal methods. WALINET provided significantly higher data quality, effectively doubling the SNR and lowering by half quantification errors of metabolites, compared to conventional methods. WALINET and LIPNET have a lipid removal factor that is 11.19 times larger than L2, and a water removal factor that is 2.86 times larger compared to HLSVD. Compared to other methods listed in the recent consensus paper tkavc2021water, WALINET and LIPNET provide lipid suppression factors similar to the crusher coil and ECLIPSE but without the need of special hardware. A potential caveat of WALINET and LIPINET is the lactate doublet at 1.3 ppm may be removed together with the lipid signal, hence the ability of to obtain lactate maps is dependent on the lactate quadruplet at 4.1 ppm. By comparison the nuisance signal removal using the union-of-space model ma2016removal has been shown to retain the lactate peak at 1.3 ppm.
Convolutional neural networks act as nonlinear functions that may model better lipid contamination in MRSI, while conventional methods bilgic2014fast, ma2016removal that assume a linear orthogonal relationship between lipids and metabolites may inadvertently remove metabolite signal when this assumption is not met. The effects of improved metabolite quantification translate into metabolite images that have better structural details. We believe that a robust self-contained efficient nuisance signal removal method that is implemented as an independent processing step is very useful and can be combined with any MRSI processing pipeline. This may offer greater flexibility compared to methods that are fully embedded with reconstruction of k-space MRSI data.
@@
At the moment, the demonstration of WALINET performance was limited to 7T ultra-high-field MRSI, which is the highest field approved for clinical use. Methods such as WALINET are highly relevant at 7T because the high SAR and non-uniform $B_{0}$ and $B_{1}+$ fields make pulse sequence-based suppression of water and lipids highly impractical for whole-brain MRSI. However, we expect that the approach and same network can be employed with additional training at lower (3T) or higher fields. We also expect that the processing time of WALINET will be similar for higher spatial resolution of MRSI, while the time for conventional methods will linearly scale with the data size. Furthermore, removal of lipid signal is extremely important for accelerated undersampled MRSI acquisitions nassirpour2018multinet, nassirpour2018compressed where aliased lipid signal can overwhelm parallel imaging and compressed sense reconstructions. Based on its performance, WALINET has great potential for being used in combination with undersampled MRSI, and we will explore this in future work. This is suggested by the generalization of WALINET from 3D Non-Cartesian sampling to 2D Cartesian sampling that we demonstrated in this paper. In addition, we showed that the same model can be used for lipid-only removal (LIPNET), which can be further combined with other processing methods. We also noticed that LIPNET provides some suppression of the water tail in the metabolite spectral range. This happens because there are lipid signals (at 5.1ppm and 4.4 ppm) close to the water peak which the network learns to remove. However, in the presence of a large residual water peak LIPNET is not sufficient and a water trained network is necessary such as WALINET.
@@ Conclusion
Efficient removal of water and lipid signals is key for proton MRSI-based quantitative metabolic imaging. Convolutional neural networks such as WALINET provide an effective approach towards this goal with superior performance compared to conventional methods. We provide WALINET, including the computation of the L2 lipid operator, as a self-contained package that can be used as a plug-in with any MRSI processing pipeline. Since WALINET does not need specialized hardware, it can be easily disseminated to automate the MRSI processing pipeline for high-throughput workflow in clinical applications. We anticipate that these aspects will lead to larger adoption and impact of MRSI for clinical applications and research.
@@ 5 Supplementary Methods
2D 1H-FID Cartesian MRSI data were acquired with unaccelerated elliptical phase-encoded. 3D 1H-FID-ECCENTRIC klauser2023eccentric was acquired with randomly positioned circular trajectories with the radius set to kmax/8 without temporal interleaving and full sampling (AF=1) of a spherical 3D k-space.
For both sequences the excitation was performed with a Shinnar-LeRoux optimized pulse Klauser2021, klauser2023eccentric having 6.5kHz bandwidth, 1ms duration and 27° excitation flip-angle. No lipid suppression was employed in the sequence, while water suppression was achieved by a four-pulses WET method Klauser2021, klauser2023eccentric.
@@
Training Data:
First, spectra for 25 common 1H metabolites were simulated using a physical model for the coupled spin systemsSmith1994, songeon2023vivo. An extensive dataset of $1.9\times 10^{6}$ metabolite spectra was simulated for a wide range of concentrations, linewidths, noise levels, and baselines. The metabolite concentrations were distributed according to a normal distribution with a standard deviation five times greater than the mean of 1 (arbitrary units), with truncation to zero to ensure positive values. The simulation included also variations in frequency offset (-150 to 150 Hz), Voigt linewidth (4 to 50 Hz), and signal-to-noise ratio (SNR from 1 to 10). Random baselines were introduced, characterized by 10 broad gaussian components.
Second, we extracted a collection of lipid spectra from voxels within the scalp region obtained from in-vivo MRSI datasets. Representative head/brain/scalp masks and B0 field maps are presented in Supplementary Figure S1. The amplitude of lipid signal from the lipid mask was varied with a scaling factor, spanning a broad interval from $10^{-2}$ to $10^{3}$. In addition, the phase of the lipid signal was varied between $\pi$ and $\pi$. The variation of the amplitude and phase of the lipid signal mimics the sinc ringing of the point spread function. The large range of amplitude and phase variation effectively augments the in-vivo lipid contamination, which makes the network robust to different k-space sampling strategies. Hence, we do not need to retrain the network when changing the acquisition. This was verified by testing WALINET on MRSI data acquired with ECCENTRIC as well as cartesian phase encoding.
To create the final training input spectra ($x_{1}$) for WALINET the simulated metabolite spectra ($m$) and experimentally-derived lipid & water spectra ($x_{2}$) were randomly combined. Note, that in the case of LIPNET the input training spectra combined simulated metabolite and experimentally-derived lipid spectra, without the water signal.
Water and lipid signals were extracted from 19 subjects, including 2 glioma patients. The MRSI data were acquired with the 3D 1H-FID-ECCENTRIC sequence described in the Methods. $10^{5}$ lipid & water spectra were extracted from each subject, resulting in $1.9\times 10^{6}$ total metabolite + lipid & water spectra used for training. Additionally, a validation dataset included 4 other subjects.
@@
For 3D ECCENTRIC the k-space sampling density was compensated based on Voronoi diagrams malik2005gridding where each k-space point is normalized by the area of its assigned Voronoi vertex. Upon the weighting of the k-space data, an inverse non-uniform discrete Fourier transform (iNUFT) was applied for each stack (kx-ky plane) within the k-space domain. Subsequently, an additional inverse fast Fourier transform (iFFT) was performed along the kz dimension to finalize the reconstruction of the MRSI data. In contrast, for 2D $^{1}H-$FID-MRSI data a iFFT is applied on the 2D k-space datasets. For both 2D and 3D MRSI a Hamming filter was applied in k-space prior to the Fourier transform. After the transformation from k space to image space, the spectra are obtained by Fourier transform of the time dimension. Coil combination was performed with ESPIRIT uecker2014espirit using sensitivity profiles computed from water un-suppressed MRSI. A correction for $B_{0}$ field inhomogeneity (Supp. Fig. S1) is computed from water un-suppressed data and applied to the coil-combined image space MRSI data, which was followed by water and lipid removal.
@@
LCModel used a basis set simulated by NMR quantum mechanics in GAMMA smith1994computer for twenty-two metabolites: phosphorylcholine (PCh), glycerophosphorylcholine (GPC), creatine (Cr), phosphocreatine (PCr), gamma-aminobutyric acid (GABA), glutamate (Glu), glutamine (Gln), glycine (Gly), glutathione (GSH), myo-inositol (Ins), N-acetylaspartate (NAA), N-acetyl aspartylglutamate (NAAG), scylloinositol (Sci), lactate (Lac), threonine (Thr), beta-glucose (bGlu), alanine (Ala), aspartate (Asp), ascorbate (Asc), serine (Ser), taurine (Tau), and 2-hydroxyglutarate (2HG) and a measured macromolecular background povavzan2015mapping. Note that during training WALINET and LIPNET learn to remove macromolecule signal, since this is present together with lipid signal in the scalp spectra used to generate training data. This was verified experimentally, as macromolecular fitting by LCModel was very close to 0 throughout the brain. The spectral fitting was done for the 1ppm-4.2ppm spectral range and the results for each voxel were used to generate metabolic images. The unsupressed water reference signal was used as quantification reference for metabolites concentrations (institutional units, I.U.) to compare metabolite levels across subjects and scanners. To assess the quality of the MRSI data and fit, linewidth (FWHM), signal-to-noise ratio (SNR), and Cramer-Rao lower bounds (CRLB) goodness of fit maps were generated.
The lipid removal factor was calculated in test subjects and in simulations. For this we integrated the absolute value of the lipid signal before and after removal: a) in the range 0.8-1.8 ppm to exclude the metabolite region (1.8-4.2 ppm) for the in-vivo data, and b) over the entire spectral range of interest 0.8-4.2 ppm in simulations, since in simulations we have the metabolite-free lipid signal over the entire range. The water removal factor was calculated by integrating the absolute water signal between 4.2-5.2 ppm, before and after the removal.
@@
Supporting Figure S1 : Examples of head/brain/scalp binary masks and the corresponding B0 field maps (specified in ppm).
@@
Supporting Figure S2 : Comparison of lipid removal on in-vivo 2D MRSI by LIPNET and L2 for three values of the regularization parameter $\beta$, the optimal value ($3.69*10^{6}$), double the optimal value, and zero for no lipid suppression. Maps are shown for two metabolites Cr+PCr, Glu and their corresponding CLRB. Spectra from several brain voxels are shown for each method, the white trace shows the measured spectrum, the red trace shows LCModel fit.
Supporting Figure S3 : Comparison of water and lipid removal to WALINET on full range spectra. Each specrtum is shown enlarged on top and in full size below. The nuisance contaminated spectra is displayed in white, HLSVD+L2 processed spectra in orange, and WALINET nuisance removed spectra in purple. The from WALINET predicted nuisance signal is overlayed in light blue color.
vol12 vol15 Simulated
Lipid Removal Factor: Brain mask Mean 25.43 45.28 235.55
20.72 25.61 17.68
Median 20.57 35.60 118.51
18.07 22.41 8.97
Lipid Removal Factor: Head mask Mean 689.23 876.14 235.55
76.81 73.59 17.68
Median 78.69 156.73 118.51
33.56 37.36 8.97
Water Removal Factor: Brain mask Mean 34.18 53.09 224.20
18.84 18.53 91.51
Median 17.19 26.40 69.81
15.54 16.43 12.54
Supporting Table S1 : Lipid and Water Removal Factor Comparison. WALINET results are shown in red, HLSVD+L2 in yellow. Lipid removal factors were computed in the head and brain, water removal factor was in the brain only. Results are shown for to subjects and on simulated data.